# An Imitative Reinforcement Learning Framework for Autonomous Dogfight

### v2

@@ Abstract.
Unmanned Combat Aerial Vehicle (UCAV) dogfight, which refers to a fight between two or more UCAVs usually at close quarters, plays a decisive role on the aerial battlefields. With the evolution of artificial intelligence, dogfighting progressively transits towards intelligent and autonomous modes. However, the development of autonomous dogfight policy learning is hindered by challenges such as weak exploration capabilities, low learning efficiency, and unrealistic simulated environments. To overcome these challenges, this paper proposes a novel imitative reinforcement learning framework, which efficiently leverages expert data while enabling autonomous exploration. The proposed framework not only enhances learning efficiency through expert imitation, but also ensures adaptability to dynamic environments via autonomous exploration with reinforcement learning. Therefore, the proposed framework can learn a successful dogfight policy of ‘pursuit-lock-launch’ for UCAVs. To support data-driven learning, we establish a dogfight environment based on the Harfang3D sandbox, where we conduct extensive experiments. The results indicate that the proposed framework excels in multistage dogfighting, and significantly outperforms state-of-the-art reinforcement learning and imitation learning methods. Thanks to the ability of imitating experts and autonomous exploration, our framework can quickly learn the critical knowledge in complex aerial combat tasks, achieving up to a $100\%$ success rate and demonstrating excellent robustness.
Reinforcement Learning, Imitation Learning, Dogfight, Air Combat
@@ 1. Introduction
In recent years, emerging artificial intelligence technologies such as deep learning have made remarkable advancements in the military field. The increasing significance of artificial intelligence applications in military operations has propelled modern warfare toward intelligent development (van2019framework, ; benaskeur2010corals, ). Unmanned combat aerial vehicles (UCAVs), as key elements in aerial combat, have become essential in modern military and defense, demonstrating outstanding combat performance in tasks such as aerial strikes, reconnaissance, electronic warfare, and target tracking. Autonomous control of UCAVs has gradually become a primary research focus (hu2022autonomous, ; chamola2021comprehensive, ; bell2017evaluating, ), with dogfight is one of the most important tasks for UCAVs, which refers to a fight between two or more UCAVs usually at close quarters. This paper mainly focuses on autonomous control policy learning for dogfight tasks between two UCAVs.
There has been substantial research on autonomous control policy learning for UCAV dogfight engagements (ruan2022autonomous, ; ramirez2018effectiveness, ; trsek2007last, ), and most existing works rely on supervised learning and offline imitation learning techniques (zhang2020maneuver, ; li2022decision, ; sandstrom2022fighter, ; kong2020uav, ). With a delicately designed dataset, these methods can learn a well-performed control policy for the UCAVs. However, the absence of online interactions can lead to the overfitting of these static datasets. In realistic combat environments, opponent policies are everchanging, and unforeseen obstacles may appear suddenly. The UCAV control policies learned with a fixed dataset can hardly handle such dynamic environments. Furthermore, the cost of constructing such a training dataset is expensive as well, and the performance of the learned policy heavily depends on the quality of the dataset. Consequently, it is of great importance to develop more adaptive learning frameworks for UCAV control policy.
@@
Recently, more researchers have investigated reinforcement learning (RL) (sutton2018reinforcement, ) techniques to achieve autonomous policy learning for UCAVs (kim2003autonomous, ; wang2022research, ; fu2022uav, ; li2023manoeuvre, ; fan2022air, ; cao2023autonomous, ). Different from supervised learning methods, RL enables the control agent for the UCAV to gradually learn policies through online interactions with the environment. Considering the dynamic property of the dogfight environment, RL is suitable for such a learning problem (piao2024discovering, ), as it supports online exploration and therefore can better adapt to dynamic changes in the environments. However, dogfight policy learning is quite a complex problem, which involves multiple control dimensions, e.g., the direction of flight, target lock, and missile launch. Due to the trial-and-error nature of RL agents, the sample efficiency of RL algorithms in such complex learning scenarios is low, leading to poor success rates of dogfight engagements.
@@
Besides the learning frameworks, the simulated learning environment is substantially important as well. Since directly conducting dogfight policy learning in real environments is impossible, an effective and high-fidelity simulator is urgently required. However, the simulated environments in previous works suffer from the oversimplified action spaces and task setting. For example, the action space in (fan2022air, ) is simplified into a discrete predefined maneuver library; (fu2022uav, ) confines the action space to a two-dimensional simulation environment; (wang2022research, ; li2023manoeuvre, ; cao2023autonomous, ) merely use part of the parameters of the UCAV flight dynamics equations as the action space. As for the task setting, (kim2003autonomous, ) solely focuses on the flight control of UCAV without launching missiles; (wu2023intelligent, ) tackles missile evasion of UCAV without flight control; (hsu2018distributed, ; choi2019deep, ; zhu2017target, ; kahn2018self, ) aim to achieve the objectives of route planning; (haarnoja2018soft, ; fu2022uav, ) aim to accomplish the pursuit subtask; (li2023manoeuvre, ) completes the locking of the opponent based on rule-based pursuing. The flight dynamics model and task setting in (berndt2004jsbsim, ) is realistic, but this simulator is implemented with C++, which is hard to adapt to the deep learning algorithms implemented with Python. To successfully shoot down the opponent in a realistic dogfight task, UCAV needs to accomplish three subtasks at least. These include swiftly pursuing the opponent through the optimal path, locking onto the target within a proper range, and launching missiles once the opponent is locked.
@@
To address the aforementioned challenges in dogfight policy learning, we introduce a novel imitative reinforcement learning framework that integrates imitation learning with autonomous exploration by the RL agent. By imitating expert data, learning efficiency is improved, while autonomous exploration by the RL agent facilitates adaptation to dynamic environments. In this work, the UCAV dogfight problem is formulated as a Markov Decision Process (MDP). To alleviate the limitations of previous dogfight simulators, we construct a realistic UCAV simulation environment based on the Harfang3D sandbox (ozbek2022harfang3d, ). The constructed learning environment features a high-fidelity action space design that enables control over UCAV directional rudders, elevators, and ailerons, making it more realistic than prior works that employed simplified dynamics equation parameters for control. Beyond the action space, the dynamic model in the constructed simulator is also of high fidelity, which is built according to the real flight dynamics models of various aircraft. As a replication of the real dogfight process, the simulated dogfight task consists of three stages, i.e., pursuit, lock, and launch, providing a more complex replication of real dogfight processes than previous single-stage works. As an evaluation, we conduct experiments in the constructed dogfight environment. The experiment results demonstrate that the training environment constructed in this paper has high fidelity, effectively simulating the UCAV dogfight process and better aligning the physical model with real-world scenarios. The proposed learning framework performs excellently in the UCAV multi-dimensional decision-making task, significantly outperforming state-of-the-art deep RL and imitation learning algorithms, achieving higher learning efficiency, more effective air combat policies, and higher success rates. The contributions of this work can be summarized as follows.
@@
  • Problem Formulation: We construct a simulated dogfight environment with high fidelity, which contains multiple stages: pursuit, lock, and launch.
@@
  • Learning Algorithm: We propose a novel imitative reinforcement learning framework for UCAV dogfight that enhances learning efficiency and adapts to complex environments.
@@
  • Empirical Improvement: The proposed framework significantly surpasses the state-of-the-art RL and imitation learning methods in the dogfight problem.
@@
This work is organized as follows. In Section 2, we introduce preliminary knowledge, and then in Section 3, we review the related works. After that, in Section 4, we formulate the dogfight problem as an MDP. In Section 5, we present the proposed framework for dogfight policy learning. Next, in Section 6, we conduct extensive experiments, and the results demonstrate that the proposed framework substantially outperforms the baselines. Finally, in Section 7, we conclude and point out the possible future directions.
@@ 2.2. Markov Decision Process
The Markov Decision Process (MDP) (garcia2013markov, ) is the theoretical foundation of RL, which provides a formal description of the learning environment of the agent. In this paper, we model the UCAV dogfight environment as an MDP. The MDP is represented as a tuple $(S,A,P,R,\gamma)$, where $S$ is the state space of the environment, $A$ is the action space of the agent, $P$ is the state transition probability, with $P(s_{t+1}|s_{t},a_{t})$ indicating the probability of transitioning from state $s_{t}$ to state $s_{t+1}$ when taking action $a_{t}$ at time $t$, $R$ is the reward function, representing the immediate reward obtained by the intelligent agent after taking action $a$, and $\gamma$ is the reward discount factor, with $\gamma\in[0,1)$ and a $\gamma$ value closer to 1 indicating that the intelligent agent focuses more on long-term rewards. Reinforcement learning aims to learn policy function $\pi:S\rightarrow A$ that maximizes discounted reward $\mathbb{E}_{P,\pi}[\sum_{t=1}^{T}\gamma^{t}R(s_{t},a_{t})]$, where $T$ is the total timestep in an episode.
@@ 3.4. Combining Reinforcement Learning and Imitation Learning
Although deep learning methods, especially deep RL, have made significant progress in UCAV dogfights, current research still faces several challenges, such as unrealistic simulation environments, oversimplified task design, low algorithmic sample efficiency, and inadequate exploration capabilities. To address these issues, we construct a realistic UCAV dogfight simulation environment based on the Harfang3D sandbox, which utilizes a UCAV control mechanism aligned with real operational procedures, and design a comprehensive dogfight task that incorporates multiple decision dimensions. Furthermore, we propose a novel imitative reinforcement learning framework that enhances learning efficiency through expert imitation. Beyond that, the proposed approach can improve the policy’s adaptability to dynamic environments through autonomous exploration, and therefore mastering complex multi-dimensional decision policies in UCAV dogfight tasks.
@@ 4. Problem Formulation
We establish a realistic simulation environment for UCAV dogfight policy learning based on the Harfang3D sandbox (ozbek2022harfang3d, ) and model the ‘pursuit-lock-launch’ tasks in a dogfight as an MDP, which is shown in Figure 2.
@@ 4.1. Task Description
In this subsection, we describe the UCAV dogfight task, which is divided into three consecutive stages: target pursuit, target lock, and missile launch. The initial position of the opponent aircraft is fixed, while the initial position of our aircraft is random, with a relative distance greater than the lock range between them. Both our aircraft and the opponent aircraft are initialized in the air. The opponent aircraft have three modes: flying forward with a constant speed, serpentine maneuvering, and circling maneuvering. In contrast, our aircraft is controlled intelligently through a neural network. If our aircraft catches up with the opponent aircraft, locks onto it, launches a missile, and finally shoots down the opponent aircraft, the mission is successful. To ensure that the agent can successfully and accurately complete this task, it needs to master at least three policies of three different stages. Successful target lock requires meeting lock conditions and maintaining them for $5$ seconds, as given below.
@@
Figure 3. An illustration of dogfight rewards. The blue aircraft denotes our aircraft, and the red one denotes the opponent aircraft.
@@ 4.2. State Space
As shown in Figure 3, within the same Cartesian coordinate system, the coordinates of our aircraft are represented as $(x_{u},y_{u},z_{u})$, and those of the opponent aircraft are denoted as $(x_{e},y_{e},z_{e})$. The relative distance vector, denoted as $\vec{D}$, is expressed as
(2) $\vec{D}=(\Delta x,\Delta y,\Delta z)=(x_{e},y_{e},z_{e})-(x_{u},y_{u},z_{u}),$
$\vec{V_{u}}$ and $\vec{V_{e}}$ are the velocity vectors of our aircraft and the opponent aircraft, respectively. The angle $q$ between the velocity vector of our aircraft and the relative distance vector is the locking angle.
@@ 4.3. Action Space
The action space consists of four types of actions: rudder, elevator, and aileron control as well as missile launch control variable. The aircraft carried only one missile, thus having one opportunity to shoot down the opponent aircraft. The rudder, elevator, and aileron control are continuous variables, and we normalize them to a range of $(-1,1)$. The missile launch control variable is discrete, with $-1$ representing no launch and $1$ representing launch. This hybrid action space, consisting of both continuous and discrete variables, presents increased challenges in dogfight policy learning, but resembles the real-world scenarios more.
@@ 4.4. Reward Function
The reward function is a critical component for applying reinforcement learning to real-world problems. In the UCAV dogfight, the operated aircraft pursues the opponent aircraft, locks its aim, and launches a missile within an extremely short time. To accomplish this task, we propose the following reward function with five terms:
@@
where $q$ is the locking angle between the velocity vector of our aircraft $\vec{V}_{u}$ and the relative distance vector $\vec{D}$.
@@ 5. Methodology
Figure 4. The proposed learning framework for dogfights.
The actor–critic framework serves as the foundation for dogfight policy learning. Inspired by the TD3 algorithm (fujimoto2018addressing, ), to mitigate the overestimation of the critic, we adopt a double Q-network architecture ($Q_{\phi_{1}}$ and $Q_{\phi_{2}}$) (van2016deep, ). Inspired by the RL objective shown in Eq.(9), the actor network $\theta_{actor}$ is trained to maximize the smaller Q value between the two Q values.
@@
where $L_{RL}$ denotes the RL loss for the actor network, $B$ represents the replay buffer, $\pi(s;\theta_{actor})$ denotes the policy network with parameter $\theta_{actor}$, and $Q(s,a)$ represents the estimated Q value by the critic network. Minimizing loss $L_{RL}$ to update the parameters of the actor network can achieve dogfight policy learning. The detailed procedure of neural network optimization includes random sampling a minibatch from the replay buffer $B$, and then optimizing $\theta_{actor}$ with stochastic gradient descent. This is standard in the deep learning domain, and it can be regarded as optimizing the expectation in Equation (9). For the rest loss functions, the same procedure is adopted.
@@
where $\alpha$ is a fixed weight to balance the magnitude $L_{RL}$ and $L_{BC}$, which is key to learning performance , and $\lambda$ is a weight to balance the two loss terms. In this work, we propose to use the loss function $(1-\lambda)L_{RL}+\alpha\lambda L_{BC}$ to update the actor network, so that we can take advantage of both RL and imitation learning. However, balancing these two learning objectives poses a significant challenge. To address this issue, we develop two mechanisms to set $\lambda$: linear $\lambda$ and adaptive $\lambda$.
@@
Algorithm 1 Imitative Reinforcement Learning Framework for Dogfight
@@
The proposed dogfight policy learning framework is outlined in Algorithm 1. First, the actor and critic networks are randomly initialized, and the target networks are synchronized. In lines 7–8, samples are collected in the simulated environment and added to the replay buffer $B$ of the first-in-first-out store manner. In lines 9–12, samples are drawn randomly from the replay buffer $B$, and the actor and critic networks are updated. The loss function for the critic network is as follows:

### v3

# An Imitative Reinforcement Learning Framework for Pursuit-Lock-Launch Missions

@@ Abstract.
Unmanned Combat Aerial Vehicle (UCAV) Within-Visual-Range (WVR) engagement, referring to a fight between two or more UCAVs at close quarters, plays a decisive role on the aerial battlefields. With the development of artificial intelligence, WVR engagement progressively advances towards intelligent and autonomous modes. However, autonomous WVR engagement policy learning is hindered by challenges such as weak exploration capabilities, low learning efficiency, and unrealistic simulated environments. To overcome these challenges, we propose a novel imitative reinforcement learning framework, which efficiently leverages expert data while enabling autonomous exploration. The proposed framework not only enhances learning efficiency through expert imitation, but also ensures adaptability to dynamic environments via autonomous exploration with reinforcement learning. Therefore, the proposed framework can learn a successful policy of ‘pursuit-lock-launch’ for UCAVs. To support data-driven learning, we establish an environment based on the Harfang3D sandbox. The extensive experiment results indicate that the proposed framework excels in this multistage task, and significantly outperforms state-of-the-art reinforcement learning and imitation learning methods. Thanks to the ability of imitating experts and autonomous exploration, our framework can quickly learn the critical knowledge in complex aerial combat tasks, achieving up to a $100\%$ success rate and demonstrating excellent robustness.
Reinforcement Learning, Imitation Learning, WVR Engagement, Air Combat
@@ 1. Introduction
In recent years, emerging artificial intelligence technologies, such as deep learning have made remarkable advancements in the military field. The increasing significance of artificial intelligence applications in military operations has propelled modern warfare toward intelligent development (van2019framework, ; benaskeur2010corals, ). Unmanned combat aerial vehicles (UCAVs), as key elements in aerial combat, have become essential in modern military and defense, demonstrating outstanding combat performance in tasks such as aerial strikes, reconnaissance, electronic warfare, and target tracking. Autonomous control of UCAVs has gradually become a primary research focus (hu2022autonomous, ; chamola2021comprehensive, ; bell2017evaluating, ), where WVR engagement is one of the most important tasks for UCAVs. WVR engagements refer to fights between two or more UCAVs usually at close quarters, which typically involve three subtasks: pursuing, locking, and launching. This paper mainly focuses on autonomous control policy learning for this pursuit-lock-launch task between two UCAVs.
@@
Recently, more researchers have investigated reinforcement learning (RL) (sutton2018reinforcement, ) techniques to achieve autonomous policy learning for UCAVs (kim2003autonomous, ; wang2022research, ; fu2022uav, ; li2023manoeuvre, ; fan2022air, ; cao2023autonomous, ). Different from supervised learning methods, RL enables the control agent for the UCAV to gradually learn policies through online interactions with the environment. Considering the dynamic property of the WVR engagement environment, RL is suitable for such a learning problem (piao2024discovering, ), as it supports online exploration and therefore can better adapt to dynamic changes in the environments. However, WVR engagement policy learning is quite a complex problem, which involves multiple control dimensions, e.g., the direction of flight, target lock, and missile launch. Due to the trial-and-error nature of RL agents, the sample efficiency of RL algorithms in such complex learning scenarios is low, leading to poor success rates of WVR engagements.
@@
Besides the learning frameworks, the simulated learning environment is substantially important as well. Since directly conducting WVR engagement policy learning in real environments is impossible, an effective and high-fidelity simulator is urgently required. However, the simulated environments in previous works suffer from the oversimplified action spaces and task setting. For example, the action space in (fan2022air, ) is simplified into a discrete predefined maneuver library; (fu2022uav, ) confines the action space to a two-dimensional simulation environment; (wang2022research, ; li2023manoeuvre, ; cao2023autonomous, ) merely use part of the parameters of the UCAV flight dynamics equations as the action space. As for the task setting, (kim2003autonomous, ) solely focuses on the flight control of UCAV without launching missiles; (wu2023intelligent, ) tackles missile evasion of UCAV without flight control; (hsu2018distributed, ; choi2019deep, ; zhu2017target, ; kahn2018self, ) aim to achieve the objectives of route planning; (haarnoja2018soft, ; fu2022uav, ) aim to accomplish the pursuit subtask; (li2023manoeuvre, ) completes the locking of the opponent based on rule-based pursuing. The flight dynamics model and task setting in JSB-sim (berndt2004jsbsim, ) are less realistic than ours, since this simulator does not integrate a missile system, burdening researchers to manually simulate missiles or utilize existing open-source repositories. To successfully shoot down the opponent in a pursuit-lock-launch task, UCAV needs to accomplish three subtasks at least: pursuiting, locking, and launching. Specifically, these include swiftly pursuing the opponent through the optimal path, locking onto the target within a proper range, and launching missiles once the opponent is locked.
@@
To address the aforementioned challenges in a pursuit-lock-launch policy learning, we introduce a novel imitative reinforcement learning framework that integrates imitation learning with autonomous exploration by the RL agent. By imitating expert data, learning efficiency is improved, while autonomous exploration by the RL agent facilitates adaptation to dynamic environments. In this work, the UCAV pursuit-lock-launch problem is formulated as a Markov Decision Process (MDP). To alleviate the limitations of previous dogfight simulators, we construct a realistic UCAV simulation environment based on the Harfang3D sandbox (ozbek2022harfang3d, ). The constructed learning environment features a high-fidelity action space design that enables control over UCAV directional rudders, elevators, and ailerons, making it more realistic than prior works that employed simplified dynamics equation parameters for control. Beyond the action space, the dynamic model in the constructed simulator is also of high fidelity, which is built according to the flight dynamics models of real aircraft. As a replication of the WVR engagement process, the simulated pursuit-lock-launch task consists of three stages, i.e., pursuit, lock, and launch, providing a more complex replication of real WVR engagement processes than previous single-stage works. As an evaluation, we conduct experiments in the constructed WVR engagement environment. The experiment results demonstrate that the training environment constructed in this paper has high fidelity, effectively simulating the UCAV WVR engagement process and better aligning the physical model with real-world scenarios. The proposed learning framework performs excellently in the UCAV multi-dimensional decision-making task, significantly outperforming state-of-the-art deep RL and imitation learning algorithms, achieving higher learning efficiency, more effective air combat policies, and higher success rates. The contributions of this work can be summarized as follows.
@@
  • Problem Formulation: We construct a simulated WVR engagement environment with high fidelity, which contains multiple stages: pursuit, lock, and launch.
@@
  • Learning Algorithm: We propose a novel imitative reinforcement learning framework for UCAV WVR engagement that enhances learning efficiency and adapts to complex environments.
@@
  • Empirical Improvement: The proposed framework significantly surpasses the state-of-the-art RL and imitation learning methods in the pursuit-lock-launch problem.
@@
This work is organized as follows. In Section 2, we introduce preliminary knowledge, and then in Section 3, we review the related works. After that, in Section 4, we formulate the pursuit-lock-launch problem as an MDP. In Section 5, we present the proposed framework for pursuit-lock-launch policy learning. Next, in Section 6, we conduct extensive experiments, and the results demonstrate that the proposed framework substantially outperforms the baselines. Finally, in Section 7, we conclude and point out the possible future directions.
@@ 2.2. Markov Decision Process
The Markov Decision Process (MDP) (garcia2013markov, ) is the theoretical foundation of RL, which provides a formal description of the learning environment of the agent. It is noteworthy that, since the intelligent agent cannot access complete details of the opponent’s information during aerial combat, modeling the scenario as a Partially Observable Markov Decision Process (POMDP) appears more theoretically justified. However, we model the UCAV WVR engagement environment as an MDP to simplify the problem in this paper. The MDP is represented as a tuple $(S,A,P,R,\gamma)$, where $S$ is the state space of the environment, $A$ is the action space of the agent, $P$ is the state transition probability, with $P(s_{t+1}|s_{t},a_{t})$ indicating the probability of transitioning from state $s_{t}$ to state $s_{t+1}$ when taking action $a_{t}$ at time $t$, $R$ is the reward function, representing the immediate reward obtained by the intelligent agent after taking action $a$, and $\gamma$ is the reward discount factor, with $\gamma\in[0,1)$ and a $\gamma$ value closer to 1 indicating that the intelligent agent focuses more on long-term rewards. Reinforcement learning aims to learn policy function $\pi:S\rightarrow A$ that maximizes discounted reward $\mathbb{E}_{P,\pi}[\sum_{t=1}^{T}\gamma^{t}R(s_{t},a_{t})]$, where $T$ is the total timestep in an episode.
@@ 3.4. Combining Reinforcement Learning and Imitation Learning
Although deep learning methods, especially deep RL, have made significant progress in UCAV dogfights, current research still faces several challenges, such as unrealistic simulation environments, oversimplified task design, low algorithmic sample efficiency, and inadequate exploration capabilities. To address these issues, we construct a realistic UCAV WVR engagement simulation environment based on the Harfang3D sandbox, which utilizes a UCAV control mechanism aligned with real operational procedures, and design a comprehensive pursuit-lock-launch task that incorporates multiple decision dimensions. Furthermore, we propose a novel imitative reinforcement learning framework that enhances learning efficiency through expert imitation. Beyond that, the proposed approach can improve the policy’s adaptability to dynamic environments through autonomous exploration, and therefore mastering complex multi-dimensional decision policies in UCAV pursuit-lock-launch tasks.
@@ 4. Problem Formulation
We establish a realistic simulation environment for UCAV WVR engagement policy learning based on the Harfang3D sandbox (ozbek2022harfang3d, ) and model the ‘pursuit-lock-launch’ tasks in a WVR engagement as an MDP, which is shown in Figure 2.
@@ 4.1. Task Description
In this subsection, we describe the UCAV WVR engagement task, which is divided into three consecutive stages: target pursuit, target lock, and missile launch. The initial position of the opponent aircraft is fixed, while the initial position of our aircraft is random, with a relative distance greater than the lock range between them. Both our aircraft and the opponent aircraft are initialized in the air. The opponent aircraft have three modes: flying forward with a constant speed, serpentine maneuvering, and circling maneuvering. In contrast, our aircraft is controlled intelligently through a neural network. If our aircraft catches up with the opponent aircraft, locks onto it, launches a missile, and finally shoots down the opponent aircraft, the mission is successful. To ensure that the agent can successfully and accurately complete this task, it needs to master at least three policies of three different stages. Successful target lock requires meeting lock conditions and maintaining them for $5$ seconds, as given below.
@@
Figure 3. An illustration of WVR engagement rewards. The blue aircraft denotes our aircraft, and the red one denotes the opponent aircraft.
@@ 4.2. State Space
As shown in Figure 3, within the same Cartesian coordinate system, the coordinates of our aircraft are represented as $(x_{u},y_{u},z_{u})$, and those of the opponent aircraft are denoted as $(x_{e},y_{e},z_{e})$. The relative position vector is denoted as $\vec{D}$$\vec{V_{u}}$ and $\vec{V_{e}}$ are the velocity vectors of our aircraft and the opponent aircraft, respectively. The angle $q$ between the velocity vector of our aircraft and the relative position vector is the locking angle, which is calculated as
(2) $q=\arccos\left(\frac{\vec{D}\cdot\vec{V_{u}}}{|\vec{D}||\vec{V_{u}}|}\right)\times\frac{180}{\pi}$
@@ 4.3. Action Space
The action space consists of four types of actions: rudder, elevator, and aileron control as well as missile launch control variable. The aircraft carried only one missile, thus having one opportunity to shoot down the opponent aircraft. The rudder, elevator, and aileron control are continuous variables, and we normalize them to a range of (1,1)(-1,1). The missile launch control variable is discrete, with 1-1 representing no launch and $1$ representing a launch. This hybrid action space, consisting of both continuous and discrete variables, presents increased challenges in pursuit-lock-launch policy learning, but resembles the real-world scenarios more.
@@ 4.4. Reward Function
The reward function is a critical component for applying reinforcement learning to real-world problems. In the UCAV pursuit-lock-launch task, the operated aircraft pursues the opponent aircraft, locks its aim, and launches a missile within an extremely short time. To accomplish this task, we propose the following reward function with five terms:
@@
where $q$ is the locking angle between the velocity vector of our aircraft $\vec{V}_{u}$ and the relative position vector $\vec{D}$.
@@ 5. Methodology
Figure 4. The proposed learning framework for WVR engagements.
The actor–critic framework serves as the foundation for pursuit-lock-launch policy learning. Inspired by the TD3 algorithm (fujimoto2018addressing, ), to mitigate the overestimation of the critic, we adopt a double Q-network architecture ($Q_{\phi_{1}}$ and $Q_{\phi_{2}}$) (van2016deep, ). Inspired by the RL objective shown in Eq.(9), the actor network $\theta_{actor}$ is trained to maximize the smaller Q value between the two Q values.
@@
where $L_{RL}$ denotes the RL loss for the actor network, $B$ represents the replay buffer, $\pi(s;\theta_{actor})$ denotes the policy network with parameter $\theta_{actor}$, and $Q(s,a)$ represents the estimated Q value by the critic network. Minimizing loss $L_{RL}$ to update the parameters of the actor network can achieve pursuit-lock-launch policy learning. The detailed procedure of neural network optimization includes random sampling a minibatch from the replay buffer $B$, and then optimizing $\theta_{actor}$ with stochastic gradient descent. This is standard in the deep learning domain, and it can be regarded as optimizing the expectation in Equation (9). For the rest loss functions, the same procedure is adopted.
@@
where $\alpha$ is a fixed weight to balance the magnitude $L_{RL}$ and $L_{BC}$, which is key to learning performance , and $\lambda$ is a weight to balance the two loss terms. We employ the most straightforward approach to set $\alpha$ : empirically determining the approximate magnitudes of both losses through several preliminary training episodes, then calculating $\alpha$ to equilibrate their scales. In this work, we propose to use the loss function $(1-\lambda)L_{RL}+\alpha\lambda L_{BC}$ to update the actor network, so that we can take advantage of both RL and imitation learning. However, balancing these two learning objectives poses a significant challenge. To address this issue, we develop two mechanisms to set $\lambda$: linear $\lambda$ and adaptive $\lambda$.
@@
Algorithm 1 Imitative Reinforcement Learning Framework for Pursuit-Lock-Launch Task
@@
The proposed pursuit-lock-launch policy learning framework is outlined in Algorithm 1. First, the actor and critic networks are randomly initialized, and the target networks are synchronized. In lines 7–8, samples are collected in the simulated environment and added to the replay buffer $B$ of the first-in-first-out store manner. In lines 9–12, samples are drawn randomly from the replay buffer $B$, and the actor and critic networks are updated. The loss function for the critic network is as follows: